OpenStudy (anonymous):

**will give medal** A membership to Movie Night Movie Club costs $10 plus $2 per movie. Y = 2x +10. Using the graph you drew for the equation, predict the amount you will spend if you rent 10 movies.

OpenStudy (the_fizicx99):

Y = 2x +10 f(x) is a fancy way of saying y, so lets use f(x) f(x) = 2x + 10 f(10) = 2(10) + 10 Multiply 2*10 and add 10 Can you continue?

OpenStudy (anonymous):

what do i do after i multiply and add ?

OpenStudy (the_fizicx99):

Well, $2 per movie you want 10 movies so multiply 2*10 and you'll have the overall cost of movies, and just add 10 for the constant 10. (membership to Movie Night Movie Club costs $10) So in other words, 2*10 = ?? ?? + 10 = ?

OpenStudy (anonymous):

30 ?

OpenStudy (the_fizicx99):

Yup, f(10) = 30

OpenStudy (anonymous):

thank you lol

OpenStudy (the_fizicx99):

yw